OpenStudy (anonymous):

graph the parabola y=3.6(x-1)^2-4

hero (hero):

Use a calculator

hero (hero):

You should find one

OpenStudy (anonymous):

tell me about it... we're not allowed to use them anyway, so it would be a waste of money!

OpenStudy (anonymous):

y=3.6(x-1)^2-4 first, look at what y=x^2 looks like. the numbers around it change the shape and location of the parabola. so -4 is the intercept, moving the center down. the -1 moves the center to the right. 3.6 measures how steep the parabola is. once you understand that, you can start to conceptually understand the graph. to start graphing it, find all the intercepts when y=0, what is x? when x=0, what is y?

OpenStudy (anonymous):

x=0.4?

OpenStudy (anonymous):

not quite. almost. plug in 0 for x. y=-0.4 so then plot (0, -.04) that is one point on your parabola. now you need more intercepts and the min.

OpenStudy (anonymous):

sorry. correction: (0,-0.4) not .04